# TrustScore: Reference-Free Evaluation of LLM Response Trustworthiness

### v1

@@ 1 Introduction
This concern is particularly salient in closed-book question-answering tasks (Roberts et al., 2020; Su et al., 2023), where the absence of contextual cues complicates the evaluation process. Unlike tasks such as reading comprehension (Yao et al., 2023), summarization (Aharoni et al., 2023), and natural language inference (Chen & Eger, 2023), where answers can be verified against provided context, assessing the trustworthiness of an LLM’s response in a closed-book setting poses a greater challenge. In such settings, LLMs rely solely on their parametric knowledge to generate answers, making it difficult to determine whether these responses truly align with their underlying knowledge due to the black-box nature of LLMs. Traditional approaches like fact-checking, which rely on evidence from external knowledge bases, may be hindered by inaccessible or insufficient information, rendering them ineffectively.
@@ 2 Proposed Method
The key component of TrustScore is a Behavioral Consistency (Trust) evaluator as illustrated in Figure 1, which assesses whether the responses provided by the LLM are consistent with its intrinsic parametric knowledge. This evaluator is designed to function independently, particularly useful in situations where external knowledge bases are not accessible, or there is not enough information available. Moreover, Trust can integrate with a Fact-Checking (Trust) module to further assess trustworthiness using an external knowledge base when one is available. This dual approach allows for a comprehensive assessment of an LLM’s response, ensuring it is evaluated both for internal consistency and external factual consistency.
2.1 Trust Stand-Alone
The Behavior Consistency Evaluator aims to determine whether the LLM consistently selects response when presented with distractors . We set a maximum check limit of times, where LLM is tasked with choosing the correct option from five choices (including response , three random distracting options, and ‘None of the above’) in each attempt. Note that the distracting options are different across different attempts. If LLM does not choose response in any of the attempts, it implies that it does not consistently provide the response to question in line with its parametric knowledge. Ideally, if we conduct an infinite number of checks and LLM consistently selects response , it implies that response is genuinely rooted in the LLM’s underlying knowledge. The behavior check process terminates when inconsistent behavior occurs (LLM does not select response ) or when it reaches the maximum check limit . If inconsistent behavior occurs, Trust return 0; otherwise, if inconsistent behavior does not occur up to the maximum check limit , Trust return 1.
@@ 2.1 Trust Stand-Alone
2.2 Integration of Trust and Trust
When external knowledge bases are available, Trust can collaborate with Trust for a thorough evaluation of responses. The role of Trust is to verify if a response to a question aligns with facts from external knowledge bases. This project does not seek to pioneer new fact-checking methods. We adopt a widely accepted approach of extracting relevant evidence from knowledge bases and assessing its support for the response through an entailment model. Trust combines scores from Trust and Trust based on criteria detailed in Appendix A.5.3. This integration is guided by several principles: (1) Responses supported by external knowledge are deemed more trustworthy than those contradicted by it. (2) A response insisted upon by LLM is more trustworthy than the response not insisted upon by LLM. (3) Factual consistency is prioritized over behavioral consistency.
@@ 3 Experiment Setup
We have created a composite dataset, MixedQA, for the purpose of analyzing the correlation between TrustScore and human evaluations. It consists of 1,000 open-ended questions randomly selected from the test sets (or development sets when test sets are unavailable) of five datasets: Natural Questions (Kwiatkowski et al., 2019), WebQuestions (Berant et al., 2013), TriviaQA (Joshi et al., 2017), HotpotQA (Yang et al., 2018), and PopQA (Mallen et al., 2022). We ignore the support context in these datasets to fit the closed-book setting. The detailed selection criteria are in the Appendix A.3.
@@
We test our TrustScore on the answer generated by FLAN-T5-XL, LLaMA-7B, and GPT-3.5-Turbo. LLMs perform the question-answering task in a few-shot setting. To fully exploit TrustScore, we prompt LLM to output response in long sentence instead of short phrase. The implementation detail including the prompts and hyper-prameters are provided in Appendix A.5.
@@
We employ the following baselines: reference-based metrics, reference-free metrics with and without evidence retrieval. Reference-based metrics include popular lexical matches like Exact Match, BLEU (Papineni et al., 2002), ROUGE-L (Lin, 2004), METEOR (Banerjee & Lavie, 2005), and pretrained evaluators such as BERTScore (Zhang et al., 2019), BARTScore (Yuan et al., 2021), and GPTScore222GPTScore includes several backbones, and we chose FLAN-T5-base in this work. (Fu et al., 2023). Reference-free metrics include BARTScore and GPTScore, both with and without evidence retrieval, for a fair comparison with our TrustScore.
@@
| Trust | 0.593 | 0.533 | 0.449 |
| Trust (Ours) | 0.613 | 0.539 | 0.484 |
@@ 4 Results
In experiments where retrieved evidence is unavailable, our Trust demonstrates the highest correlation with human judgments, outperforming other reference-free metrics. Remarkably, Trust also outperforms some reference-based metrics when evaluating responses from GPT-3.5. This is particularly significant given that, in many real-world scenarios, relevant evidence is unavailable or insufficient, making reference-based or retrieval-based metrics inapplicable. In such cases, our Trust can effectively assess response trustworthiness. Comparing the Trust scores across different LLMs, we observe that the Trust score for LLaMA is lower than those for FLAN-T5 and GPT3.5. This difference could be attributed to LLaMA being a base model, while FLAN-T5 is fine-tuned with instructions and GPT-3.5 is fine-tuned with both instructions and human feedback. These fine-tuning processes may enhance the behavioral consistency of the latter models.
Integration of Trust and Trust
In experiments where retrieved evidence is available, Trust demonstrates human correlations that surpass those of existing reference-free metrics and even outperform most reference-based metrics. We note that the Trust score consistently outperforms Trust across various LLMs. This observation substantiates our assertion that Trust is not only an effective reference-free metric on its own but can also complement Trust to achieve better performance in a reference-based setting.
@@ 6 Conclusion
We introduce TrustScore, a novel reference-free evaluation framework that assesses the trustworthiness of an LLM response from (1) behavior consistency of the LLM in making coherent choices among its response and other distractor, and (2) the factuality of their generated response. Our experiments, conducted on our newly curated MixedQA dataset, demonstrate that the two components complement each other effectively within the TrustScore framework. TrustScore exhibits a strong correlation with human judgments, surpassing existing reference-free metrics, and achieves results comparable to reference-based metrics.
@@ A.3 Data Selection Criteria for MixedQA
Input to Language Models: We input the filtered questions, devoid of any contextual information, to three language models: FLAN-T5-XL, LLaMA-7B, and GPT-3.5-Turbo.
@@ A.5.1 Details of the Substitution Algorithm
To generate high-quality distractors, we propose a substitution algorithm that produces a dictionary of substitute candidates , where keys represent words to be substituted and the values are the corresponding potential substitute words. As shown in Algorithm1, the substitution algorithm involves three functions to search for substitute candidates for a given entity or word: find_dbpedia_ents, find_candidates_by_embedding, and find_candidate_by_pos. The find_dbpedia_ents function retrieves DBpedia entities with the same DBpedia types as the entities to be substituted. We employ the DBpedia Spotlight for Spacy to detect DBpedia entities in response and obtain their corresponding DBpedia types. For example, the DBpedia types for an entity like ”Kobe” might include ”DBpedia:Person,” ”DBpedia:Athlete,” and ”DBpedia:BasketballPlayer.” Subsequently, we use SPARQL to search for DBpedia entities that satisfy these types, starting with entities that match all types and gradually lowering the number of required matching types if necessary. The find_candidates_by_embedding function identifies the most similar words by computing cosine similarity between the mean of the projection weight vectors of the words to be substituted and the vectors for each word in a pretrained word2vec model (word2vec-google-news-300). The find_candidate_by_pos function randomly selects words with the same POS tag from the word2vec-google-news-300 corpus.

### v2

@@ 1 Introduction
This concern is particularly salient in closed-book question-answering tasks (Roberts et al., 2020; Su et al., 2023), where the absence of contextual cues complicates the evaluation process. Unlike tasks such as reading comprehension (Yao et al., 2023), summarization (Aharoni et al., 2023), and natural language inference (Chen & Eger, 2023), where answers can be verified against provided context, assessing the trustworthiness of an LLM’s response in a closed-book setting poses a greater challenge. In such settings, LLMs rely solely on their parametric knowledge (Pan et al., 2023) to generate answers, making it difficult to determine whether these responses truly align with their underlying knowledge due to the black-box nature of LLMs. Traditional approaches like fact-checking, which rely on evidence from external knowledge bases (Pan et al., 2017a; b), may be hindered by inaccessible or insufficient information, rendering them ineffectively.
@@ 2 Proposed Method
The key component of TrustScore is a Behavioral Consistency (TrustBC) evaluator as illustrated in Figure 1, which assesses whether the responses provided by the LLM are consistent with its intrinsic parametric knowledge. This evaluator is designed to function independently, particularly useful in situations where external knowledge bases are not accessible, or there is not enough information available. Moreover, TrustBC can integrate with a Fact-Checking (TrustFC) module to further assess trustworthiness using an external knowledge base when one is available. This dual approach allows for a comprehensive assessment of an LLM’s response, ensuring it is evaluated both for internal consistency and external factual consistency.
2.1 TrustBC Stand-Alone
The Behavior Consistency Evaluator aims to determine whether the LLM consistently selects response when presented with distractors . We set a maximum check limit of times, where LLM is tasked with choosing the correct option from five choices (including response , three random distracting options, and ‘None of the above’) in each attempt. Note that the distracting options are different across different attempts. If LLM does not choose response in any of the attempts, it implies that it does not consistently provide the response to question in line with its parametric knowledge. Ideally, if we conduct an infinite number of checks and LLM consistently selects response , it implies that response is genuinely rooted in the LLM’s underlying knowledge. The behavior check process terminates when inconsistent behavior occurs (LLM does not select response ) or when it reaches the maximum check limit . If inconsistent behavior occurs, TrustBC return 0; otherwise, if inconsistent behavior does not occur up to the maximum check limit , TrustBC return 1.
@@ 2.1 TrustBC Stand-Alone
2.2 Integration of TrustBC and TrustFC
When external knowledge bases are available, TrustBC can collaborate with TrustFC for a thorough evaluation of responses. The role of TrustFC is to verify if a response to a question aligns with facts from external knowledge bases. This project does not seek to pioneer new fact-checking methods. We adopt a widely accepted approach of extracting relevant evidence from knowledge bases and assessing its support for the response through an entailment model. TrustOV combines scores from TrustBC and TrustFC based on criteria detailed in Appendix A.5.3. This integration is guided by several principles: (1) Responses supported by external knowledge are deemed more trustworthy than those contradicted by it. (2) A response insisted upon by LLM is more trustworthy than the response not insisted upon by LLM. (3) Factual consistency is prioritized over behavioral consistency.
@@ 3 Experiment Setup
We have created a composite dataset, MixedQA, for the purpose of analyzing the correlation between TrustScore and human evaluations. It consists of 1,000 open-ended questions randomly selected from the test sets (or development sets when test sets are unavailable) of five datasets: Natural Questions (Kwiatkowski et al., 2019), WebQuestions (Berant et al., 2013), TriviaQA (Joshi et al., 2017), HotpotQA (Yang et al., 2018), and PopQA (Mallen et al., 2022). We ignore the support context in these datasets to fit the closed-book setting. The detailed selection criteria are in the Appendix A.3. The dataset and code are public available 222https://github.com/dannalily/TrustScore.
@@
We test our TrustScore on the answer generated by FLAN-T5-XXL, LLaMA-7B, and GPT-3.5-Turbo. LLMs perform the question-answering task in a few-shot setting. To fully exploit TrustScore, we prompt LLM to output response in long sentence instead of short phrase. The implementation detail including the prompts and hyper-prameters are provided in Appendix A.5.
@@
We employ the following baselines: reference-based metrics, reference-free metrics with and without evidence retrieval. Reference-based metrics include popular lexical matches like Exact Match, BLEU (Papineni et al., 2002), ROUGE-L (Lin, 2004), METEOR (Banerjee & Lavie, 2005), and pretrained evaluators such as BERTScore (Zhang et al., 2019), BARTScore (Yuan et al., 2021), and GPTScore333GPTScore includes several backbones, and we chose FLAN-T5-base in this work. (Fu et al., 2023). Reference-free metrics include BARTScore and GPTScore, both with and without evidence retrieval, for a fair comparison with our TrustScore.
@@
| TrustFC | 0.593 | 0.533 | 0.449 |
| TrustOV (Ours) | 0.613 | 0.539 | 0.484 |
@@ 4 Results
In experiments where retrieved evidence is unavailable, our TrustBC demonstrates the highest correlation with human judgments, outperforming other reference-free metrics. Remarkably, TrustBC also outperforms some reference-based metrics when evaluating responses from GPT-3.5. This is particularly significant given that, in many real-world scenarios, relevant evidence is unavailable or insufficient, making reference-based or retrieval-based metrics inapplicable. In such cases, our TrustBC can effectively assess response trustworthiness. Comparing the TrustBC scores across different LLMs, we observe that the TrustBC score for LLaMA is lower than those for FLAN-T5 and GPT3.5. This difference could be attributed to LLaMA being a base model, while FLAN-T5 is fine-tuned with instructions and GPT-3.5 is fine-tuned with both instructions and human feedback. These fine-tuning processes may enhance the behavioral consistency of the latter models.
Integration of TrustBC and TrustFC
In experiments where retrieved evidence is available, TrustOV demonstrates human correlations that surpass those of existing reference-free metrics and even outperform most reference-based metrics. We note that the TrustOV score consistently outperforms TrustFC across various LLMs. This observation substantiates our assertion that TrustBC is not only an effective reference-free metric on its own but can also complement TrustFC to achieve better performance in a reference-based setting.
@@ 6 Conclusion
We introduce TrustScore, a novel reference-free evaluation framework that assesses the trustworthiness of LLM responses. Our experiments, conducted on our newly curated MixedQA dataset, demonstrate effectiveness of the TrustScore framework. TrustScore exhibits a strong correlation with human judgments, surpassing existing reference-free metrics, and achieves results comparable to reference-based metrics. As for future work, we might look into how to combine TrustScore with model editing methods (Han et al., 2023).
Acknowledgments
This work is supported by Huawei’s Dean’s Funding (C-00006589) and the UKRI Centre for Doctoral Training in Natural Language Processing, funded by the UKRI (grant EP/S022481/1).
@@ A.3 Data Selection Criteria for MixedQA
Input to Language Models: We input the filtered questions, devoid of any contextual information, to three language models: FLAN-T5-XXL, LLaMA-7B, and GPT-3.5-Turbo.
@@ A.5.1 Details of the Substitution Algorithm
To generate high-quality distractors, we propose a substitution algorithm that produces a dictionary of substitute candidates , where keys represent words to be substituted and the values are the corresponding potential substitute words. As shown in Algorithm1, the substitution algorithm involves three functions to search for substitute candidates for a given entity or word: find_dbpedia_ents, find_candidates_by_embedding, and find_candidate_by_pos. The find_dbpedia_ents function retrieves DBpedia entities with the same DBpedia types (Hu et al., 2022; 2023) as the entities to be substituted. We employ the DBpedia Spotlight for Spacy to detect DBpedia entities in response and obtain their corresponding DBpedia types. For example, the DBpedia types for an entity like ”Kobe” might include ”DBpedia:Person,” ”DBpedia:Athlete,” and ”DBpedia:BasketballPlayer.” Subsequently, we use SPARQL to search for DBpedia entities that satisfy these types, starting with entities that match all types and gradually lowering the number of required matching types if necessary. The find_candidates_by_embedding function identifies the most similar words by computing cosine similarity between the mean of the projection weight vectors of the words to be substituted and the vectors for each word in a pretrained word2vec model (word2vec-google-news-300). The find_candidate_by_pos function randomly selects words with the same POS tag from the word2vec-google-news-300 corpus.